# Discontinuous Galerkin discretization of conservative dynamical low-rank approximation schemes for the Vlasov–Poisson equation

### v1

@@ 2.4 Rank-adaptive unconventional integrator
The rank $r$ in the DLRA model needs to be large enough to ensure accuracy of the solution, but still (if possible) small enough to keep the model efficient. Since determining a suitable value of $r$ is necessary, the missing rank-adaptivity of the projector splitting integrator poses a practical disadvantage. The rank-adaptive integrator proposed in [4] addresses this issue by first computing the K- and L-step independently, and then using basis augmentation in the S-step. The procedure is as follows.
@@
    (Basis augmentation) Let $r^{\prime}=2r$, $\{\tilde{X}^{1}_{1},\dots,\tilde{X}^{1}_{r^{\prime}}\}$ be an orthonormal basis (in $L_{2}(\Omega^{(x)})$) for the augmented set $[\bm{X}^{0},\bm{K}^{1}]$, and $\{\tilde{V}^{1}_{1},\dots,\tilde{V}^{1}_{r^{\prime}}\}$ an orthonormal basis (in $L_{2}(\Omega^{(v)};\omega)$) for $[\bm{V}^{0},\bm{L}^{1}]$. Then we can represent $f(t_{0})$ as
@@
    (S-step) Perform an S-step for the matrix $\tilde{S}$. However, instead of the “affine” modified S-step from (28), where only the part $\widehat{S}$ is evolved, here we need a “full” S-step as in the usual unconventional integrator. Also note that in contrast to the projector splitting integrator, the S-step in the unconventional integrator is performed “forward”, that is, using the formulation as in (11) with “$+$” and test functions of the form
@@
The idea of basis augmentation is actually quite general. Instead of $\bm{K}^{1}$ and $\bm{L}^{1}$ one could use other or even more functions for basis enrichment. Notably, it has been emphasized in [10] that in order to ensure conservation laws in the S-step, the new spatial basis should also contain all gradients $\partial_{x_{s}}\nabla X^{0}_{i}$ of the initial basis functions collected as $\nabla\bm{X}^{0}$. For the conservative unconventional integrator one therefore takes a new orthonormal basis $\tilde{X}_{1}^{1},\tilde{X}_{2}^{1},\dots$ of the augmented system $[\bm{X}^{0},\nabla\bm{X}^{0},\bm{K}^{1}]$ instead of just $[\bm{X}^{0},\bm{K}^{1}]$.
The above procedure would increase the rank in every time step and needs to be combined with rank truncation. Following [10] this is achieved as follows.
@@ 3.1 Discontinuous Galerkin discretization for Friedrichs’ systems
We first recall the discretization of Friedrichs’ systems using the DG approach. Our formulation closely follows [6, Chapter 8]. Other references include [16] and [12]. In this work, we restrict ourselves to the case of periodic domains or to problems without boundary (i.e. on $\mathbb{R}^{d}$). In the latter case, however, we assume that the computational domain can be taken sufficiently large to fully include the support of the solution in its interior (up to machine precision) and hence can also be treated as a periodic domain numerically. We note that in principle, DG discretizations also work for domains with boundary, but this would require some more careful considerations, which will be left for future work.
@@
Let $\mathcal{F}_{h}^{(y)}$ be the set of all $(d-1)$-dimensional faces of elements in $\mathcal{T}_{h}^{(y)}$. Since we assume that the domain is periodic, all faces are in fact internal faces. For each $e\in\mathcal{F}_{h}^{(y)}$ we define a fixed unit normal vector $\bm{n}^{(e)}$ (the direction can be arbitrary). Taking into account that the domain is periodic, every boundary $\partial T$ of an element essentially decomposes into shared faces with its neighboring elements. For every face $e\in\mathcal{F}_{h}^{(y)}$, there exist two elements $T^{+}$ and $T^{-}$ such that $e\subset T^{+}\cap T^{-}$. We use the convention that $T^{+}$ lies in the direction of $\bm{n}^{(e)}$. Given a vector field $\bm{U}$ on $\Omega^{(y)}$, one then defines the average and jump of $\bm{U}$,
@@
$\langle\hat{\mathrm{d}}_{y_{s}}U_{h},V_{h}\rangle_{\mathcal{T}_{h}^{(y)}}\coloneqq\sum_{T\in\mathcal{T}_{h}^{(y)}}\int_{T}\partial_{y_{s}}U_{h}\,V_{h}\,\mathrm{d}\bm{y}-\sum_{e\in\mathcal{F}_{h}^{(y)}}\int_{e}n_{s}^{(e)}\,[U_{h}]_{e}\,\{V_{h}\}_{e}\,\mathrm{d}S.$ (33)
@@
$\langle\hat{\mathrm{d}}_{y_{s}}U_{h},V_{h}\rangle_{\mathcal{T}_{h}^{(y)}}=-\sum_{T\in\mathcal{T}_{h}^{(y)}}\int_{T}U_{h}\partial_{y_{s}}V_{h}\,\mathrm{d}\bm{y}+\sum_{e\in\mathcal{F}_{h}^{(y)}}\int_{e}n_{s}^{(e)}\{U_{h}\}_{e}\,[V_{h}]_{e}\,\mathrm{d}S.$
@@
$\langle\hat{\mathrm{d}}_{y_{s}}U_{h},V_{h}\rangle_{\mathcal{T}_{h}^{(y)}}=-\sum_{T\in\mathcal{T}_{h}^{(y)}}\int_{T}U_{h}\partial_{y_{s}}V_{h}\,\mathrm{d}\bm{y}+\sum_{e\in\mathcal{F}_{h}^{(y)}}\int_{e}n_{s}^{(e)}[U_{h}\,V_{h}]_{e}\,\mathrm{d}S-\sum_{e\in\mathcal{F}_{h}^{(y)}}\int_{e}n_{s}^{(e)}\,[U_{h}]_{e}\,\{V_{h}\}_{e}\,\mathrm{d}S$
@@ 3.2 Application of DG discretization to DLRA
$\mathsf{A}^{(x,s)}=\Big{[}\big{(}v_{s}V_{j}^{n},V_{i}^{n}\big{)}_{v,\omega}\Big{]}_{ij},\qquad\mathsf{B}_{h}^{(x,s)}=\Big{[}\langle\hat{\mathrm{d}}_{v_{s}}(\omega V_{j}^{n}),V_{i}^{n}\rangle_{\mathcal{T}_{h}^{(v)}}\Big{]}_{ij},$ (34)
@@
$\mathsf{A}^{(v,s)}=\Big{[}\big{(}-E_{s}X_{j}^{n},X_{i}^{n}\big{)}_{x}\Big{]}_{ij},\qquad\mathsf{B}_{h}^{(v,s)}=\Big{[}\langle\hat{\mathrm{d}}_{x_{s}}X_{j}^{n},X_{i}^{n}\rangle_{\mathcal{T}_{h}^{(x)}}\Big{]}_{ij}.$
@@
Finally, for the S-step we use the unconventional integrator with an augmented basis. We define orthonormal bases $\tilde{\bm{X}}^{n+1}\in(\mathcal{V}_{h,p}^{(x)})^{r(2+d)}$ and $\tilde{\bm{V}}^{n+1}\in(\mathcal{V}_{h,p}^{(v)})^{2r-m}$ for the augmented matrices $[\bm{X}^{n},\hat{\bm{\Phi}}^{n},\bm{K}^{n+1}]$ and $[\bm{V}^{n},\bm{L}^{n+1}]$, respectively. Here, $\hat{\bm{\Phi}}^{n}$ contains all Riesz representatives of partial derivatives $\hat{\Phi}_{X^{n}_{i},s}$, and $\bm{L}^{n+1}$ contains the new basis functions $L^{n+1}_{p}$ from the L-step. The orthonormal basis $\tilde{\bm{V}}^{n+1}$ is always chosen such that the first $m$ functions remain the fixed $U_{a}$.
@@
with $k$ running from $1,\dots,r(2+d)$ and $\ell$ from $1,\dots,2r-m$. It follows that the state at the previous ($n$-th) step can be expressed in the new basis as
$\sum_{ij}S_{ij}^{n}X_{i}^{n}\otimes V_{j}^{n}=\sum_{k\ell}\left(\sum_{ij}M_{ki}S_{ij}^{n}N_{\ell j}\right)\tilde{X}_{k}^{n+1}\otimes\tilde{V}_{\ell}^{n+1}.$ (36)
@@
$f(t)=\omega\sum_{k\ell}S_{k\ell}(t)\tilde{X}^{n+1}_{k}\otimes\tilde{V}^{n+1}_{\ell},$
@@
$\begin{split}S_{k\ell}^{n+1}=M_{ki}\cdot S_{ij}^{n}\cdot N_{\ell j}-\tau\sum_{ijs}\Big{[}&\langle\hat{\mathrm{d}}_{x_{s}}X_{i}^{n},\tilde{X}_{k}^{n+1}\rangle_{\mathcal{T}_{h}^{(x)}}\,S_{ij}^{n}\,(\tilde{V}_{\ell}^{n+1},v_{s}V_{j}^{n})_{v,\omega}\\ &+(\tilde{X}_{k}^{n+1},-E_{s}X_{i}^{n})_{x}\,S_{ij}^{n}\,\langle\hat{\mathrm{d}}_{v_{s}}(\omega V_{j}^{n}),\tilde{V}_{\ell}^{n+1}\rangle_{\mathcal{T}_{h}^{(v)}}\Big{]},\end{split}$ (37)
@@ 3.3 Conservation of mass and momentum
In this section, we establish the mass and momentum conservation properties of the unconventional integrator applied to the DG discretization. Specifically, following [10], we prove that the numerical scheme satisfies discrete continuity relations consistent with the analytic continuity equations. As a weight function $\omega(\bm{v})$, we use the Gaussian weight (12), although generalizations should be possible.
@@
Consider now $f=\omega\sum_{ij}S_{ij}X_{i}\otimes V_{j}$ with fixed functions $V_{1}=1/\|1\|_{v,\omega}$ and $V_{1+s}=v_{s}/\|v_{s}\|_{v,\omega}$. For the Gaussian weight, these fixed functions are already pairwise orthonormal in the weighted inner product on $\Omega^{(v)}=\mathbb{R}^{d}$. As before, we assume that the whole system of $V_{j}$’s is orthonormal in this inner product. The mass density and momentum density of $f$ can then be expressed as
$\rho$ $=\sum_{ij}X_{i}S_{ij}(\omega V_{j},1)_{v}=\|1\|_{v,\omega}\sum_{ij}X_{i}S_{ij}(V_{j},V_{1})_{v,\omega}=\|1\|_{v,\omega}\sum_{i}X_{i}S_{i1},$ (40)
$j_{s}$ $=\sum_{ij}X_{i}S_{ij}(\omega V_{j},v_{s})_{v}=\|v_{s}\|_{v,\omega}\sum_{ij}X_{i}S_{ij}(V_{j},V_{1+s})_{v,\omega}=\|v_{s}\|_{v,\omega}\sum_{i}X_{i}S_{i,1+s}$ (41)
for $s\in\{1,\ldots,d\}$. Note that we could also write $\rho=\|1\|_{v,\omega}K_{1}$ and $j_{s}=\|v_{s}\|_{v,\omega}K_{1+s}$. Total mass and momentum conservation hence means that the integrals of $K_{1},\dots,K_{1+d}$ remain constant.
@@
In the numerical scheme from section 3.2 discrete quantities $\rho^{n}$ and $j^{n}_{s}$ are computed for $n=0,1,2\dots,$ from basis functions $X_{i}^{n}\in\mathcal{V}_{h,p}^{(x)}$ and coefficient matrices $S^{n}$. The goal is to show local continuity relations for the discrete quantities. The key is the discrete integration by parts formula in Lemma 3.1.
@@
$\frac{1}{\tau}(\rho^{n+1}-\rho^{n},\varphi)_{x}-\sum_{T\in\mathcal{T}_{h}^{(x)}}\int_{T}\bm{j}^{n}\cdot\nabla\varphi\,\mathrm{d}x+\sum_{e\in\mathcal{F}_{h}^{(x)}}\int_{e}\big{(}\bm{n}^{(e)}\cdot\{\bm{j}^{n}\}_{e}\big{)}\,[\varphi]_{e}\,\mathrm{d}S=0$ (42)
for all $\varphi\in\mathcal{V}_{h,p}^{(x)}$. This corresponds to a DG discretization of the continuity equation (38) using central fluxes and explicit Euler time stepping as in (32). Hence, the discrete local continuity equation holds for the mass density, and the total mass $\int_{\Omega^{(x)}}\rho^{n}\,\mathrm{d}\bm{x}$ is conserved.
@@
By combining (40) with (36) and (37), the updated mass density can be expressed as
$\rho^{n+1}=\|1\|_{v,\omega}\sum_{k}\tilde{X}^{n+1}_{k}S^{n+1}_{k1}=\rho^{n}+\tau\Delta\rho$ (43)

### v2

@@ 2.4 Rank-adaptive unconventional integrator
The rank $r$ in the DLRA model needs to be large enough to ensure accuracy of the solution, but still (if possible) small enough to keep the model efficient. Since determining a suitable value of $r$ is necessary, the missing rank-adaptivity of the projector splitting integrator poses a practical disadvantage. The rank-adaptive integrator proposed in [5] addresses this issue by first computing the K- and L-step independently, and then using basis augmentation in the S-step. It has been adopted to the conservative setup with fixed velocity functions in [12, 10]. The procedure is as follows.
@@
    (Basis augmentation) Let $r^{\prime}=2r$, $\{\tilde{X}^{1}_{1},\dots,\tilde{X}^{1}_{r^{\prime}}\}$ be an orthonormal basis (in $L_{2}(\Omega^{(x)})$) for the augmented set $\bm{X}^{0},\bm{K}^{1}]$, and $\{\tilde{V}^{1}_{1},\dots,\tilde{V}^{1}_{r^{\prime}}\}$ an orthonormal basis (in $L_{2}(\Omega^{(v)};\omega)$) for $\bm{V}^{0},\bm{L}^{1}]$. Then we can represent $f(t_{0})$ as
@@
    (S-step) Perform an S-step for the matrix $\tilde{S}$. However, instead of the “affine” modified S-step from (28), where only the part $\widehat{S}$ is evolved, here we need a “full” S-step as in the usual unconventional integrator. Also note that in contrast to the projector splitting integrator, the S-step in the unconventional integrator is performed “forward”, that is, using the formulation as in (10) with “++” and test functions of the form
@@
The idea of basis augmentation is actually quite general. Instead of $\bm{K}^{1}$ and $\bm{L}^{1}$ one could use other or even more functions for basis enrichment. The above procedure would increase the rank in every time step and needs to be combined with rank truncation. Following [12] this is achieved as follows.
@@ 3.1 Discontinuous Galerkin discretization for Friedrichs’ systems
We first recall the discretization of Friedrichs’ systems using the DG approach. Our formulation closely follows [7, Chapter 8]. Other references include [18] and [14]. In this work, we restrict ourselves to the case of periodic domains or to problems without boundary (i.e. on $\mathbb{R}^{d}$). In the latter case, however, we assume that the computational domain can be taken sufficiently large to fully include the support of the solution in its interior (up to machine precision) and hence can also be treated as a periodic domain numerically. We note that in principle, DG discretizations also work for domains with boundary, but this would require some more careful considerations, which will be left for future work.
@@
Let $\mathcal{F}_{h}^{(y)}$ be the set of all (d1)(d-1)-dimensional faces of elements in $\mathcal{T}_{h}^{(y)}$. Since we assume that the domain is periodic, all faces are in fact internal faces. For each $e\in\mathcal{F}_{h}^{(y)}$ we define a fixed unit normal vector $\bm{n}^{(e)}$ (the direction can be arbitrary). Taking into account that the domain is periodic, every boundary $\partial T$ of an element essentially decomposes into shared faces with its neighboring elements. For every face $e\in\mathcal{F}_{h}^{(y)}$, there exist two elements $T^{+}$ and $T^{-}$ such that $e\subset T^{+}\cap T^{-}$. We use the convention that $T^{+}$ lies in the direction of $\bm{n}^{(e)}$. Given a vector field $\bm{U}$ on $\Omega^{(y)}$, one then defines the average and jump of $\bm{U}$,
@@
$\big{(}\hat{\mathrm{d}}_{y_{s}}U_{h},V_{h}\big{)}_{y}\coloneqq\sum_{T\in\mathcal{T}_{h}^{(y)}}\int_{T}\partial_{y_{s}}U_{h}\,V_{h}\,\mathrm{d}\bm{y}-\sum_{e\in\mathcal{F}_{h}^{(y)}}\int_{e}n_{s}^{(e)}\,[U_{h}]_{e}\,\{V_{h}\}_{e}\,\mathrm{d}S.$ (33)
@@
$\big{(}\hat{\mathrm{d}}_{y_{s}}U_{h},V_{h}\big{)}_{y}=-\sum_{T\in\mathcal{T}_{h}^{(y)}}\int_{T}U_{h}\partial_{y_{s}}V_{h}\,\mathrm{d}\bm{y}+\sum_{e\in\mathcal{F}_{h}^{(y)}}\int_{e}n_{s}^{(e)}\{U_{h}\}_{e}\,[V_{h}]_{e}\,\mathrm{d}S.$
@@
$\big{(}\hat{\mathrm{d}}_{y_{s}}U_{h},V_{h}\big{)}_{y}=-\sum_{T\in\mathcal{T}_{h}^{(y)}}\int_{T}U_{h}\partial_{y_{s}}V_{h}\,\mathrm{d}\bm{y}+\sum_{e\in\mathcal{F}_{h}^{(y)}}\int_{e}n_{s}^{(e)}[U_{h}\,V_{h}]_{e}\,\mathrm{d}S-\sum_{e\in\mathcal{F}_{h}^{(y)}}\int_{e}n_{s}^{(e)}\,[U_{h}]_{e}\,\{V_{h}\}_{e}\,\mathrm{d}S$
@@ 3.2 Application of DG discretization to DLRA
$\mathsf{A}^{(x,s)}=\Big{[}\big{(}v_{s}V_{j}^{n},V_{i}^{n}\big{)}_{v,\omega}\Big{]}_{ij},\qquad\mathsf{B}_{h}^{(x,s)}=\Big{[}\big{(}\hat{\mathrm{d}}_{v_{s}}(\omega V_{j}^{n}),V_{i}^{n}\big{)}_{v}\Big{]}_{ij},$
@@
$\mathsf{A}^{(v,s)}=\Big{[}\big{(}-E_{s}X_{j}^{n},X_{i}^{n}\big{)}_{x}\Big{]}_{ij},\qquad\mathsf{B}_{h}^{(v,s)}=\Big{[}\big{(}\hat{\mathrm{d}}_{x_{s}}X_{j}^{n},X_{i}^{n}\big{)}_{x}\Big{]}_{ij}.$
@@
Finally, for the S-step we use the unconventional integrator with an augmented basis. We define orthonormal bases $\tilde{\bm{X}}^{n+1}\in(\mathcal{V}_{h,p}^{(x)})^{2r}$ and $\tilde{\bm{V}}^{n+1}\in(\mathcal{V}_{h,p}^{(v)})^{2r-m}$ for the augmented matrices $\bm{X}^{n},\bm{K}^{n+1}]$ and $\bm{V}^{n},\bm{L}^{n+1}]$, respectively. Here, $\bm{L}^{n+1}$ contains the new basis functions $L^{n+1}_{p}$ from the L-step. The orthonormal basis $\tilde{\bm{V}}^{n+1}$ is always chosen with respect to the $(\cdot,\cdot)_{v,\omega}$ inner product and such that the first $m$ functions remain the fixed $U_{a}$.
@@
with $k$ running from $1,\dots,2r$ and $\ell$ from $1,\dots,2r-m$. It follows that the state at the previous ($n$-th) step can be expressed in the new basis as
$\sum_{ij}S_{ij}^{n}X_{i}^{n}\otimes V_{j}^{n}=\sum_{k\ell}\tilde{S}^{n}_{k\ell}\tilde{X}_{k}^{n+1}\otimes\tilde{V}_{\ell}^{n+1},\quad\tilde{S}^{n}_{k\ell}=\sum_{ij}M_{ki}S_{ij}^{n}N_{\ell j}.$ (35)
@@
$f(t)=\omega\sum_{k\ell}\tilde{S}_{k\ell}(t)\tilde{X}^{n+1}_{k}\otimes\tilde{V}^{n+1}_{\ell},$
@@
$\begin{split}\tilde{S}_{k\ell}^{n+1}=\tilde{S}_{k\ell}^{n}-\tau\sum_{ijs}\Big{[}&\big{(}\hat{\mathrm{d}}_{x_{s}}X_{i}^{n},\tilde{X}_{k}^{n+1}\big{)}_{x}\,S_{ij}^{n}\,(\tilde{V}_{\ell}^{n+1},v_{s}V_{j}^{n})_{v,\omega}\\ &+(\tilde{X}_{k}^{n+1},-E_{s}X_{i}^{n})_{x}\,S_{ij}^{n}\,\big{(}\hat{\mathrm{d}}_{v_{s}}(\omega V_{j}^{n}),\tilde{V}_{\ell}^{n+1}\big{)}_{v}\Big{]},\end{split}$ (36)
@@ 3.3 Conservation of mass and momentum
In this section, we establish the mass and momentum conservation properties of the unconventional integrator applied to the DG discretization when using central fluxes. Specifically, following [10, 12], we prove that the numerical scheme satisfies discrete continuity relations consistent with the analytic continuity equations. As a weight function $\omega(\bm{v})$, we use the Gaussian weight (11), although generalizations are possible.
@@
In the numerical scheme from section 3.2, instead of $\rho$ and $j_{s}$, discrete quantities $\rho^{n}$ and $j^{n}_{s}$ are computed for $n=0,1,2\dots,$ from the discrete bases $\bm{X}^{n},\bm{V}^{n}$ and coefficient matrices $S^{n}$. The goal is to establish local continuity relations for the discrete quantities. For this we adapt the approach recently presented in [10].
Lemma 3.2.
For $f^{n}=\omega\sum_{ij}S^{n}_{ij}X^{n}_{i}\otimes V^{n}_{j}$ define
$\hat{F}(f^{n}):=\sum_{s}-v_{s}\,\hat{\mathrm{d}}_{x_{s}}f^{n}+E^{n}_{s}\,\hat{\mathrm{d}}_{v_{s}}f^{n}.$
Then, using the central flux, i.e., $\alpha=1$ in (31), the K-step (34) can be written as
$(K^{n+1}_{j},\psi)_{x}=(K^{n}_{j},\psi)_{x}+\tau\big{(}\hat{F}(f^{n}),\psi\otimes V^{n}_{j}\big{)}_{xv},\quad j=1,\ldots,r,$ (39)
for all $\psi\in\mathcal{V}^{(x)}_{h,p}$. We also have
$\tilde{S}^{n+1}_{k\ell}=\tilde{S}^{n}_{k\ell}+\tau\big{(}\hat{F}(f^{n}),\tilde{X}^{n+1}_{k}\otimes\tilde{V}^{n+1}_{\ell}\big{)}_{xv}$ (40)
for the S-step (36).
Proof.
For the first equation, recall that $f^{n}=\omega\sum_{j}K^{n}_{j}\otimes V^{n}_{j}$, which gives
$\big{(}\hat{F}(f^{n}),\psi\otimes V^{n}_{k}\big{)}_{xv}$ $=\sum_{s}\sum_{j}\big{(}-v_{s}\omega V_{j}^{n},V^{n}_{k}\big{)}_{v}\cdot\big{(}\hat{\mathrm{d}}_{x_{s}}K^{n}_{j},\psi\big{)}_{x}+\big{(}\hat{\mathrm{d}}_{v_{s}}(\omega V_{j}^{n}),V^{n}_{k}\big{)}_{v}\cdot\big{(}E^{n}_{s}K^{n}_{j},\psi\big{)}_{x}$
$=\sum_{s}\sum_{j}-\mathsf{A}^{(x,s)}_{kj}\,\big{(}\hat{\mathrm{d}}_{x_{s}}K^{n}_{j},\psi\big{)}_{x}-\mathsf{B}^{(x,s)}_{h,kj}\,\big{(}E^{n}_{s}K^{n}_{j},\psi\big{)}_{x}\,.$
Now, Lemma 3.1 implies
$\big{(}\hat{\mathrm{d}}_{x_{s}}K^{n}_{j},\psi\big{)}_{x}=-\sum_{T\in\mathcal{T}_{h}^{(x)}}\int_{T}K^{n}_{j}\partial_{x_{s}}\psi\,\mathrm{d}\bm{x}+\sum_{e\in\mathcal{F}_{h}^{(x)}}\int_{e}n_{s}^{(e)}\{K^{n}_{j}\}_{e}\,[\psi]_{e}\,\mathrm{d}S.$
Plugging this into the previous equation yields shows that (39) is just (34) in component form for central fluxes with $\alpha=1$.
The second equation is immediate from (36) and the representation of $f^{n}$. ∎
Next note that both the discrete density and momentum density can be written in the form $\varphi_{U}^{n}=\big{(}f^{n},U\big{)}_{v}$ for appropriate functions $U$: for the density $\rho$ we take $U=1$ whereas for $j_{s}$ we take $U=v_{s}$. By construction, all these functions are elements of the subspace $\operatorname{span}(U_{a})$ spanned by the fixed velocity functions, see section 2.2. We now derive a discrete equation of motion for $\varphi_{U}^{n}$ which eventually will eventually lead to discrete versions of the continuity equations (37) and (38).
Lemma 3.3.
Let $U\in\operatorname{span}(U_{a})$ and define
$\varphi_{U}^{n}=\big{(}f^{n},U\big{)}_{v}\in\mathcal{V}^{(x)}_{h,p}$.
Then, these quantities satisfy the update equation
$\big{(}\varphi_{U}^{n+1},\psi\big{)}_{x}=\big{(}\varphi_{U}^{n},\psi\big{)}_{x}+\tau\Big{(}\big{(}\hat{F}(f^{n}),U\big{)}_{v},\psi\Big{)}_{x}\quad\text{for all }\psi\in V^{(x)}_{h,p}$ (41)
when the algorithm described in section 3.2 is applied with central fluxes.
Proof.
Let $P$ denote the orthogonal projection of $\big{(}\hat{F}(f^{n}),U)_{v}$ onto $\mathcal{V}^{(x)}_{h,p}$. We first show that $P\in\operatorname{span}(\tilde{\bm{X}}^{n+1})$. Recall that
$U=\sum_{j}\big{(}V^{n}_{j},U\big{)}_{v,\omega}V^{n}_{j}$ (42)
since, by construction, $U$ belongs to $\operatorname{span}({\bm{V}}^{n})$ and in the algorithm it is always ensured that the $V^{n}_{j}$ form an orthonormal system in the weighted $(\cdot,\cdot)_{v,\omega}$ inner product. Therefore, it holds that
$\Big{(}\big{(}\hat{F}(f^{n}),U)_{v},\psi\Big{)}_{x}$ $=\Big{(}\sum_{j}\big{(}\hat{F}(f^{n}),V^{n}_{j}\big{)}_{v}\,\big{(}V^{n}_{j},U\big{)}_{v,\omega},\psi\Big{)}_{x}=\sum_{j}\big{(}\hat{F}(f^{n}),\psi\otimes V^{n}_{j}\big{)}_{xv}\,\big{(}V^{n}_{j},U\big{)}_{v,\omega}$
$=\Big{(}\sum_{j}\frac{1}{\tau}\big{[}K^{n+1}_{j}-K^{n}_{j}\big{]}\,\big{(}V^{n}_{j},U\big{)}_{v,\omega},\psi\Big{)}_{x}$
for all $\psi\in\mathcal{V}^{(x)}_{h,p}$, where we used (39). This shows
$P=\sum_{j}\frac{1}{\tau}\big{[}K^{n+1}_{j}-K^{n}_{j}\big{]}\,\big{(}V^{n}_{j},U\big{)}_{v,\omega}\in\operatorname{span}(\tilde{\bm{X}}^{n+1})$
since $K^{n}_{j},K^{n+1}_{j}\in\operatorname{span}(\tilde{\bm{X}}^{n+1})$. As a result, since the $\tilde{X}_{i}^{n+1}$ are an orthonormal system by construction, we can write
$P=\sum_{i}\Big{(}\tilde{X}^{n+1}_{i},\big{(}\hat{F}(f^{n}),U\big{)}_{v}\Big{)}_{x}\,\tilde{X}^{n+1}_{i}$
and hence it holds that
$\Big{(}\big{(}\hat{F}(f^{n}),U)_{v},\psi\Big{)}_{x}=\Big{(}\sum_{i}\Big{(}\tilde{X}^{n+1}_{i},\big{(}\hat{F}(f^{n}),U\big{)}_{v}\Big{)}_{x}\,\tilde{X}^{n+1}_{i},\psi\Big{)}_{x}$ (43)
for all $\psi\in\mathcal{V}^{(x)}_{h,p}$. We can now prove (41). Using (40), we have
$\big{(}\varphi^{n+1}_{U},\psi\big{)}_{x}$ $=\Big{(}\big{(}\omega\sum_{ij}\tilde{S}^{n+1}_{ij}\,\tilde{X}^{n+1}_{i}\otimes\tilde{V}^{n+1}_{j},\,U\big{)}_{v},\psi\Big{)}_{x}$
$=\Big{(}\sum_{ij}\Big{(}\omega\Big{[}\tilde{S}^{n}_{ij}+\tau\big{(}\hat{F}(f^{n}),\tilde{X}^{n+1}_{i}\otimes\tilde{V}^{n+1}_{j}\big{)}_{xv}\Big{]}\tilde{X}^{n+1}_{i}\otimes\tilde{V}^{n+1}_{j},\,U\Big{)}_{v},\psi\Big{)}_{x}$
$=\big{(}\varphi^{n}_{U},\psi\big{)}_{x}+\tau\Big{(}\sum_{ij}\Big{(}\omega\big{(}\hat{F}(f^{n}),\tilde{X}^{n+1}_{i}\otimes\tilde{V}^{n+1}_{j}\big{)}_{xv}\tilde{X}^{n+1}_{i}\otimes\tilde{V}^{n+1}_{j},\,U\Big{)}_{v},\psi\Big{)}_{x}.$
Due to (42) the last expression equals
$\big{(}\varphi^{n+1}_{U},\psi\big{)}_{x}=\big{(}\varphi^{n}_{U},\psi\big{)}_{x}+\tau\Big{(}\sum_{i}\tilde{X}^{n+1}_{i}\big{(}\hat{F}(f^{n}),\tilde{X}^{n+1}_{i}\otimes U\big{)}_{xv},\psi\Big{)}_{x},$
which in light of (43) shows (41). Finally, this relation is preserved if the conservative rank truncation (29) from [12] is applied. ∎
Based on Lemma 3.3 we now establish the discrete continuity equations.
@@
$\frac{1}{\tau}(\rho^{n+1}-\rho^{n},\psi)_{x}-\sum_{T\in\mathcal{T}_{h}^{(x)}}\int_{T}\bm{j}^{n}\cdot\nabla\psi\,\mathrm{d}x+\sum_{e\in\mathcal{F}_{h}^{(x)}}\int_{e}\big{(}\bm{n}^{(e)}\cdot\{\bm{j}^{n}\}_{e}\big{)}\,[\psi]_{e}\,\mathrm{d}S=0$ (44)
for all $\psi\in\mathcal{V}_{h,p}^{(x)}$. This corresponds to a DG discretization of the continuity equation (37) using central fluxes and explicit Euler time stepping as in (32). Hence, the discrete local continuity equation holds for the mass density, and the total mass $\int_{\Omega^{(x)}}\rho^{n}\,\mathrm{d}\bm{x}$ is conserved.
@@
Letting $U=1$, so that $\rho=\varphi_{U}$, and applying Lemma 3.3, we obtain
$\rho^{n+1}=\rho^{n}+\tau\Big{(}\big{(}\hat{F}(f^{n}),1\big{)}_{v},\psi\Big{)}_{x}\,$,